The perimeter of a rectangle is 120 inches. The width is 12 in less than the length. What are the dimensions of the rectangle
Answer by stanbon(75887)

The perimeter of a rectangle is 120 inches. The width is 12 in less than the length. What are the dimensions of the rectangle
--------
P = 2(L + W)
120 = 2(L + (L-12))
----
60 = 2L - 12
2L = 72
L = 36 inches
------
Width = L-12 = 24 inches
-----------------
Cheers,
Stan H.
=================